OpenStudy (anonymous):

f(x) = x^3*sin(x+sqrt(x+cos2x) I get f'(x) = 3x^2*sin(x+sqrt(x+cos2x))+x^3*cos(x+(cos2x)^(-1/2))*(1+(-sin(2x)))*2.... Can someone look at this and tell me if it's right? Test tomorrow need to make sure I understand this!

OpenStudy (anonymous):

f(x) = x^3*sin(x+sqrt(x+cos2x)) sorry

OpenStudy (anonymous):

Let's see... \[\small \begin{align*} f(x)&=x^3\sin(x+\sqrt{x+\cos2x})\\\\ f'(x)&=3x^2\sin(x+\sqrt{x+\cos2x})+x^3\cos(x+\sqrt{x+\cos2x})\frac{d}{dx}[x+\sqrt{x+\cos2x}]\\\\ &=\cdots+x^3\cos(x+\sqrt{x+\cos2x})\left(1+\frac{1}{2\sqrt{x+\cos2x}}\frac{d}{dx}[x+\cos2x]\right)\\\\ &=\cdots+x^3\cos(x+\sqrt{x+\cos2x})\left(1+\frac{1}{2\sqrt{x+\cos2x}}\left(1-\sin2x\frac{d}{dx}[2x]\right)\right)\\\\ &=\cdots+x^3\cos(x+\sqrt{x+\cos2x})\left(1+\frac{1}{2\sqrt{x+\cos2x}}\left(1-2\sin2x\right)\right) \end{align*}\]

OpenStudy (anonymous):

Looks like my answer is the same to me, right? Except I didn't include the whole chain rule in the same (...), but I'll note that!

OpenStudy (anonymous):

Perhaps you have it right on paper, but what you typed is \[3x^2\sin(x+\sqrt{x+\cos2x})+2x^3\cos(x+(\cos2x)^{-1/2})(1-\sin2x)\] which is not correct.

OpenStudy (anonymous):

Yeah I screwed up typing it in a couple places I think, thanks for the help!